OpenStudy (anonymous):

If f(x) = x − 4 and g(x) = 1 over 2x − 2, find f[g(−6)].

OpenStudy (anonymous):

what the... @lgbasallote, why did you tell her I could fix this? you meanie -__- :P

OpenStudy (lgbasallote):

she asked for you >:)) hint @rebeccaskell94 evaluate g(6) first..substitute 6 to every x in g(x)

OpenStudy (lgbasallote):

i mean substitute -6*

OpenStudy (anonymous):

would the answer be -3?

OpenStudy (lgbasallote):

@rebeccaskell94 can take it from here @kcute_face_bhunnie-_- no :)

OpenStudy (lgbasallote):

oh wait...1 over..didnt see that lemme re-solve

OpenStudy (anonymous):

wait i got it itss 9?

OpenStudy (lgbasallote):

g(x) = \(\frac{1}{2x -2}\) right???

OpenStudy (anonymous):

these are my choices 9 3 −9 −3

OpenStudy (lgbasallote):

or is it g(x) = \(\frac{1}{2x} -2\)

hero (hero):

1 over 2x - 2 means 1 is on top and 2x - 2 is on the bottom

OpenStudy (anonymous):

Captain Obvious has arrived ;)

OpenStudy (lgbasallote):

i know o.O im not getting an answer in the choices

hero (hero):

Thanks for the compliment rebecca :( How friendly you are

OpenStudy (lgbasallote):

the only one that makes sense is g(x) = \(frac{1}{2} x - 2\)

OpenStudy (lgbasallote):

\(\frac{1}{2} x - 2\)

OpenStudy (lgbasallote):

@rebeccaskell94 can answer it now..

hero (hero):

Okay, yeah, that's probably it then. I'm going to go back and hide in my mouse hole and not disturb anyone else. Especially @rebeccaskell94

OpenStudy (anonymous):

that's the only one that makes sense so g(-6) = (1/2)(-6) - 2 g(-6) = -3 - 2 g(-6) = -5 substitute to x of f(x) f(g(-6)) = -5 - 4 = -9 Soorry @hero....you take things personally a lot though :( Thanks @lgbasallote

OpenStudy (lgbasallote):

hahaha fail =))) your spaces makes it confusing :p g(-6) = (1/2)(-6) - 2 g(-6) = -3 - 2 g(-6) = -5 substitute to x of f(x) f(g(-6)) = -5 - 4 = -9

hero (hero):

I don't go around calling people captain obvious though.

OpenStudy (anonymous):

close enough ;) @lgba

OpenStudy (anonymous):

thanks :)